OpenStudy (anonymous):

Use the distributive property to simplify the expression below: 2x-8 over 2. x − 8 x − 4 2x − 4 x + 2

OpenStudy (anonymous):

Simplify the following expression. 2x − 6y + 2z − 5x − 3y + 8z Use the distributive property to simplify the expression below: 1/2 (g − 8) Simplify the following expression. 2x − 8y + 3x^2 + 7y − 12x -Use the distributive property to simplify the expression below: −6(2u − 1) -Use the distributive property to simplify the expression below: −2(5d − 1) + 4d -Use the distributive property to simplify the expression below: 5t − 6(1 + t) -Evaluate: 4 − 49 ÷ 7 + 3^2 - 3x-9/-3

OpenStudy (anonymous):

it's smthin like a(b+c) =a*b + a*c similarly (b+c)/a=b/a+c/a (2x-8)/2=2x/2-8/2

OpenStudy (anonymous):